OpenStudy (anonymous):

I need help with solving systems of equations using substitutions for algebra. Not looking for answers or anything. Just need help learning it :)

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin4.htm try this

OpenStudy (amistre64):

in a system of equation, the x and y values are consistent for each equation. By "solving" for a variable in one equation; you can substitute that value into the other equation and effectively reducing it to one equation in one unknown. Solve for the unknown and you have the means to solve for the missing variable.

OpenStudy (anonymous):

Thanks guys! This helped alot! Can anyone provide an example?

OpenStudy (mathstudent55):

Example: Solve the following system of equations using the substitution method. 2x + y = 10 x - y = 2 1. First we solve one equation for one variabe. Let's solve the second equation for x: x - y = 2 Add y to both sides: x = y + 2 2. Now we substitute what x is equal to in the other equation: The first equation is 2x + y = 10. We replace x with y + 2: 2(y + 2) + y = 10. Since we have one equation with one variable, we can solve for y. Distribute the 2 on the left side: 2y + 4 + y = 10 Collect terms on left side: 3y + 4 = 10 Subtract 4 from both sides: 3y = 6 Divide both sides by 3: y = 2. Now that we know what y is equal to, we substitute the value of into either one of the original equations. Let's use the second equation: x - y = 2, but y = 2, so x - 2 = 2 Add 2 to both sides: x = 4 Solution: x = 2, y = 4